A disc of moment of inertia of a body about s given axis 15kgm^2 rotates constantly about the centre with an angular velocity of 20rads^-1. find its rotational energy and its angular momentum about the centre
1
Expert's answer
2019-05-24T10:39:58-0400

Denote "I" - moment of inertia and "\omega" -angular velocity.

Then rotational energy is


"E=\frac{I\omega^2}{2}=\frac{15\times20^2}{2}=3000 J"

and angular momentum is


"L=I\omega=20\times 15=300 \frac{kg\times m}{s}"